# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 2.6 LMC X--3
Swift J1727.816131727.816131727.8-16131727.8 - 1613 is a recently discovered BH-XRB transient, first detected by Swift/BAT on August 24, 2023. Immediate monitoring of the source with MAXI/GSC in the $2-20$ keV energy range revealed an ‘unusual’ peak in X-ray flux, increasing from $150$ mCrab to $7$ Crab within a few days of detection Nakajima et al. (2023); Nandi et al. (2024). Radio counterpart of the source, consistent with its optical position, was detected at $5.25$ GHz with VLA and $5$ GHz with ATA, showing a rise in radio flux from $\sim 18$ to $107$ mJy six days after discovery. Recent measurements estimate the source distance as $\sim 2.7$ kpc, with a spin of $\sim 0.98$ and an inclination of $\sim 40^{\circ}$ Mata Sánchez et al. (2024); Peng et al. (2024a). Further, Nandi et al. (2024) reported the detection of evolving type-C QPOs in hard X-rays (up to $\sim 100$ keV) using AstroSat, likely associated with the inner Comptonizing corona. However, NICER observations suggest a vertically elongated, jet-like corona based on the evolution of type-C QPOs Liao et al. (2024).
2.8 GX 339--4
@@
The transient BH-XRB GX $339-4$ was discovered by the OSO77-7- 7 mission in $1973$ Markert et al. (1973) and typically undergoes outbursts every $2-3$ years. Since its discovery, it has remained one of the most extensively studied BH-XRBs due to its rich spectro-temporal variability Méndez & van der Klis (1997); Corbel et al. (2003); Zdziarski et al. (2004); Belloni et al. (2005); Nandi et al. (2012); Aneesha et al. (2024). Detailed timing studies across multiple outbursts have identified distinct temporal features, including type-A, type-B, and type-C QPOs, which are closely associated with the canonical spectral states exhibited by the source (Belloni et al., 2005; Nandi et al., 2012; Mondal et al., 2023; Zhang et al., 2024; Aneesha et al., 2024, and references therein). The black hole mass is constrained to be within $8.28-11.89$ $\rm M_{\odot}$ Sreehari et al. (2019a), while the source distance is estimated as $8.4\pm 0.9$ kpc Parker et al. (2016). Recent studies suggest the source inclination angle between $30^{\circ}-34^{\circ}$, though spin measurements remain highly model-dependent, yielding values ranging from negative to moderate positive spins Zdziarski et al. (2024).
2.9 Swift J151857.0--572147
@@
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 is a newly discovered Galactic BH-XRB transient, first detected by Swift/BAT during its $2024$ outburst Kennea et al. (2024). Follow-up radio observations with ATCA at $5.5$ and $9$ GHz confirmed the presence of bright radio flares Carotenuto & Russell (2024), generally observed for the BH-XRB transients during the transition from hard to soft spectral states. Subsequent observations with NICER and Insight-HXMT constrained the source distance, inclination, and spin parameter as $\sim 5.8$ kpc, $\sim 2.1^{\circ}$, and $\sim 0.84$, respectively Peng et al. (2024b). Additionally, coherent type-C QPOs were detected with Insight-HXMT in the intermediate state of the outburst, which have been linked to the shock instability model of transonic accretion flow around black holes Chatterjee et al. (2024).
2.10 IGR J17091--3624
@@
IGR J17091362417091362417091-362417091 - 3624 is a galactic BH-XRB transient system, discovered by INTEGRAL/IBIS in $2003$ (Kuulkers et al., 2003). The source exhibits periodic outbursts with a quiescent period of typically four years between the successive outbursts (Capitanio et al., 2009; Krimm et al., 2011; Miller et al., 2016; Katoch et al., 2021). Interestingly, IGR J17091362417091362417091-362417091 - 3624 exhibits several exotic X-ray variability signatures, including the well known “heartbeat” class and therefore, often termed as the twin of GRS $1915+105$ (Capitanio et al., 2012; Radhika et al., 2018; Katoch et al., 2021; Wang et al., 2024; Shui et al., 2024). Despite multiple outbursts over the last two decades, the mass, inclination, and distance of the source are not well constrained due to the null-detection of its optical counterpart. However, several indirect measurements predict the source parameters. Using reflection spectroscopy, the inclination of the source is estimated as $\sim 30^{\circ}-45^{\circ}$ (Xu et al., 2017; Wang et al., 2018). In addition, spectral modeling suggests a black hole of mass $10-12.3$ $M_{\odot}$ likely to be present at the center of IGR J17091362417091362417091-362417091 - 3624 (Iyer et al., 2015; Radhika et al., 2018). Based on the luminosity estimates, the distance of the source is predicted to lie within $11-17$ kpc (Rodriguez et al., 2011).
2.11 MAXI J1744--294
@@
MAXI J174429417442941744-2941744 - 294, located near the Galactic center, is the most recently discovered X-ray transient by MAXI/GSC on January 2, 2025 at a flux level of $\sim 133$ mCrab (Kudo et al., 2025). The energy spectra of the source, obtained from follow-up NuSTAR observations, are found to be well described with the model comprising a thermal disc component of temperature $\sim 0.7$ keV, a power-law component of photon index $\sim 2.3$ and iron line profile at $\sim 6.6$ keV (Mandel et al., 2025). Based on these initial findings, the source is inferred to be a BH-XRB transient observed in the HSS (Mandel et al., 2025). Notably, subsequent NICER monitoring of MAXI J174429417442941744-2941744 - 294 shows an even steeper photon index of $\sim$3 (Jaisawal et al., 2025). Assuming a distance of $8$ kpc, the luminosity of the source is estimated to be $\sim 1.5\times 10^{37}$ erg $\rm s^{-1}$ in $2-10$ keV energy range (Mandel et al., 2025). Further, Grollimund et al. (2025) reported the detection of the radio counterpart of the source at $\sim 1.3$ GHz with MeerKAT.
@@ 3 Data reduction
IXPE, NASA’s first space-based low-energy X-ray polarimetry mission, was launched on December 9, 2021 (Weisskopf et al., 2022). It comprises three identical gas-pixel detector units (DUs) designed to measure the polarimetric properties of astrophysical sources using the principle of photo-electron tracking in the $2-8$ keV energy band. For each detected photon, the Stokes parameters (I, Q, U) are determined based on the azimuthal angle of the electric field vector, which is reconstructed by analyzing the photo-electron track within the detector plane. The additive nature of the Stokes parameters enables the summation of individual photon contributions to obtain the resultant I, Q, and U values, which are then used to calculate the polarization degree (PD) and polarization angle (PA).
@@
For the polarimetric analysis, we utilize the cleaned and calibrated level-2 event files from the three DUs of IXPE in the $2-8$ keV energy range. Data reduction is performed using the latest IXPEOBSSIMv31.0.3 software (Baldini et al., 2022), following standard procedures outlined in Kislat et al. 2015; Strohmayer 2017; Kushwaha et al. 2023b; Majumder et al. 2024a, b. The source and background regions are defined as a $60$ arcsec circular region centered at the source coordinates and an annular region with radii between $180$ and $240$ arcsec, respectively (Kushwaha et al., 2023b; Majumder et al., 2024a). The XPSELECT task is used to extract source and background events from these regions. We use XPBIN task with various algorithms, such as PCUBE, PHA1, PHA1Q and PHA1U, to generate the necessary data products for model-dependent polarization as well as spectro-polarimetric studies, respectively. The latest response files provided by the software team [^3] https://ixpeobssim.readthedocs.io/en/latest/ are used during the fitting of Stokes spectra. The IXPE light curves of $2000$ s time bin are extracted from the event files of the DUs using XSELECT.
@@
NICER (Gendreau et al., 2016) is the X-ray mission onboard the International Space Station (ISS), sensitive in the soft X-ray energy band of $0.2-12$ keV and capable of excellent observations for in-depth spectro-temporal studies. The NICER/XTI data is reduced using the standard data extraction software (NICERDASv13444https://heasarc.gsfc.nasa.gov/docs/nicer/nicer_analysis.html) integrated within HEASOFT V6.34555https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/RelNotes_6.34.html with the appropriate calibration database. We use the tool nicerl2 with standard filtering criteria to generate the clean event files. The spectral products are generated using the task nicerl3-spect with background model 3c50 (Remillard et al., 2022). The NICER light curves of desired time resolutions are generated using nicerl3-lc task of the software routine.
@@
We analyze the NuSTAR (Harrison et al., 2013) data using the dedicated software NUSTARDAS [^6] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/ available in HEASOFT V6.34. We use the task nupipeline to generate the cleaned event files from both FPMA and FPMB instrument onboard NuSTAR. A circular region of $60$ arcsec radii at the source position and away from it is considered to extract the source and background spectra, instrument response and ancillary files, respectively, using the task nuproduct. All the NuSTAR spectra are grouped with $25$ counts per bin to obtain better statistics during the spectral fitting.
@@
AstroSat (Agrawal, 2006) data are extracted using the latest reduction software LaxpcSoftv3.4.4777http://www.tifr.res.in/~astrosat_laxpc/LaxpcSoft.html.The standard procedures of AstroSat/LAXPC data extraction are performed following Antia et al. 2017, 2021; Antia et al. 2022; Sreehari et al. 2019b, 2020; Majumder et al. 2022. We use LAXPC20 data for generating the light curves in different energy ranges for the respective sources except Swift J1727.816131727.816131727.8-16131727.8 - 1613, for which LAXPC10 data in low gain mode was available for analysis (see Nandi et al., 2024, for details). Note that, level-1 data corresponding to single events from all layers of the detector are considered in each observation.
@@
We utilize on-demand [^8] http://maxi.riken.jp/mxondem/ processed MAXI/GSC (Matsuoka et al., 2009) data and scaled map transient analysis data [^9] https://swift.gsfc.nasa.gov/results/transients/ of Swift/BAT (Krimm et al., 2013) for the individual sources in different energy ranges. Note that, MAXI/GSC count rates in $2-20$ keV energy range are converted into the flux unit (Crab) following the relation $1$ Crab $=3.8$ photons $\rm cm^{-2}$ $\rm s^{-1}$ as suggested by the instrument team [^10] http://maxi.riken.jp/top/readme.html. Similarly, for Swift/BAT data in the $15-50$ keV energy range, we obtain the corresponding photon flux [^11] https://swift.gsfc.nasa.gov/results/transients/ in units of mCrab, where $1$ mCrab $=0.000220$ counts $\rm cm^{-2}$ $\rm s^{-1}$.
@@ 4 Analysis and Results
In this section, we present the comprehensive analyses, modeling, and observational results of the BH-XRBs under consideration, observed with IXPE. This study incorporates quasi-simultaneous observations from NICER, NuSTAR and AstroSat to enable wide-band spectro-temporal investigations. Furthermore, we explore the polarimetric properties with IXPE and examine the spectro-polarimetric correlations for these sources.
@@ 4.1 Temporal Analysis
Figure 2: Results of temporal (top panel) and spectral (bottom panel) analyses of ten BH-XRBs in different spectral states obtained using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations. Top panels: PDS obtained from NICER ($0.5-10$ keV) and/or NuSTAR/AstroSat observations in different energy ranges are depicted for the respective sources in various epochs. The power spectrum of GX $339-4$ from NICER observation is rescaled by a factor of $5$. Bottom panels: Energy spectra of the sources from NICER and/or NuSTAR data are shown during the respective epochs. PDS and energy spectra in LHS, HIMS, SIMS, HSS and SPL states are presented using red, blue, magenta, orange and green colors, respectively. PDS from NuSTAR and AstroSat data are plotted using circle and triangle, irrespective of the spectral states. See the text for details.
@@ 4.1.1 Multi-mission Monitoring
We study the variability properties of each source using multi-mission data during the epochs of IXPE observations. In Fig. 1, we present the count rate obtained with MAXI/GSC ($2-20$ keV) and Swift/BAT ($15-50$ keV) monitoring in panels (a) and (b) along with the variation of hardness ratio ($HR$) from MAXI/GSC in panel (c), respectively, for all sources under consideration. Note that, flux values associated with the count rates in soft (hard) energy bands with MAXI/GSC (Swift/BAT) are also mentioned in units of mCrab in the $y$-axis (right side) of the respective panels. Further, radio flux densities obtained from different observational campaigns, such as RATAN (Veledina et al., 2024a; Ingram et al., 2024), E-MERLIN (Williams-Baldwin et al., 2023), VLITE (Peters et al., 2023) and AMI-LA (Steiner et al., 2024) near the IXPE epochs, are shown in panel (d) for the respective sources based on their availability. We observe that all sources show marginal variation in both MAXI/GSC and Swift/BAT count rates over the entire IXPE exposure for a given epoch. Hence, in Fig. 1a-c, we present the average values of the respective quantities over the entire IXPE exposure of each epoch.
@@
Interestingly, for Swift J1727.816131727.816131727.8-16131727.8 - 1613, we observe that both MAXI/GSC and Swift/BAT flux decrease as its outburst progresses. The $HR$ also sharply decreases and anti-correlates with the accompanied radio flux density (see Fig. 1c-d). The variation of both low and high energy fluxes along with $HR$ in presence of radio ejection suggest that the source evolves through spectral state transitions. In particular, we notice exceptionally bright ($\gtrsim 3$ Crab) intermediate states ($HR\sim 0.1-0.5$, E1-E5) followed by relatively faint ($\sim 30$ mCrab) softer states ($HR\sim 0.04-0.1$, E6-E7) before the appearance of dimmed ($\sim 100$ mCrab) hard state ($HR\sim 0.3$, E8) (see Table 2). Moreover, these variability features resemble the ‘canonical’ state transition as HIMS $\rightarrow$ SIMS $\rightarrow$ HSS $\rightarrow$ LHS (decay), commonly exhibited by BH-XRB transients (see Nandi et al., 2012, for details). Similarly, for Cyg X11-1- 1, we observe a marginal increase in the MAXI/GSC count rate and a corresponding decrease in the Swift/BAT count rate during epochs E3 to E7. In the subsequent observations (E8 to E11), the MAXI/GSC count rate decreases while the Swift/BAT count rate rises. In addition, $HR$ shows a significant drop from $0.4$ to $0.1$ followed by an increase in values exceeding $0.4$. This trend possibly indicates the spectral state transitions between LHS (E1-E2, E8-E11) and HSS (E3-E7, E12), respectively (see Table 1).
@@
However, only marginal variations are seen in both low- and high-energy count rates for several sources, such as LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J174429417442941744-2941744 - 294. For all of them, the hardness ratio remains low ($HR\lesssim 0.1$) and these sources are also detected at flux levels below approximately $100$ mCrab. These possibly suggest that the sources are likely in softer spectral states. In contrast, Cyg X33-3- 3 exhibits a clear sequence of state transitions, evolving from the LHS to the HSS, then to the SIMS, and back to HSS. This evolution is accompanied by moderate changes in count rates and hardness ratio, and is likely associated with strong radio flares ($\sim 100$ mJy) (see Table 1, Fig. 1, and Veledina et al. 2024a). For 4U $1630-47$, HSS (Kushwaha et al., 2023b) and SPL (Rodriguez Cavero et al., 2023) states are observed with reasonable variations in the count rates and $HR$. It is worth noting that IGR $17091-3624$ is observed in the LHS at $\sim 1$ Crab MAXI/GSC flux with $HR\sim 0.4$, very similar to the LHS observations of Cyg X11-1- 1.

### v2

@@ 2 Source Selection and observations
$\bullet$ Swift J1727.8-1613: This is a recently discovered BH-XRB transient, first detected by Swift/BAT on August 24, 2023. Immediate monitoring of the source with MAXI/GSC in the 2202-20 keV energy range revealed an ‘unusual’ peak in X-ray flux, increasing from $150$ mCrab to $7$ Crab within a few days of detection Nakajima et al. (2023); Nandi et al. (2024). Recent measurements estimate the source distance as $\sim 2.7$ kpc, with a spin of $\sim 0.98$ and an inclination of $\sim 40^{\circ}$ Mata Sánchez et al. (2024); Peng et al. (2024a). X-ray polarimetric constraints with IXPE in various spectral states of Swift J1727.816131727.8-1613 also aim to unveil the disc-corona geometry of the source (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
@@
$\bullet$ GX 339-4: It was discovered by the OSO7-7 mission in $1973$ Markert et al. (1973) and typically undergoes outbursts every 232-3 years. The black hole mass is constrained to be within 8.2811.898.28-11.89 $\rm M_{\odot}$ Sreehari et al. (2019a), while the source distance is estimated as $8.4\pm 0.9$ kpc Parker et al. (2016). Recent studies suggest the source inclination angle between $30^{\circ}-34^{\circ}$, though spin measurements remain highly model-dependent, yielding values ranging from negative to moderate positive spins Zdziarski et al. (2024). Further, Mastroserio et al. (2025) reported the alignment of X-ray polarization direction with that of radio jet orientation in GX 3394339-4.
@@
$\bullet$ Swift J151857.0-572147: This is a newly discovered Galactic BH-XRB transient, first detected by Swift/BAT during its $2024$ outburst Kennea et al. (2024). Subsequent observations with NICER and Insight-HXMT constrained the source distance, inclination, and spin parameter as $\sim 5.8$ kpc, $\sim 2.1^{\circ}$, and $\sim 0.84$, respectively Peng et al. (2024b). Additionally, coherent type-C QPOs were detected with Insight-HXMT in the intermediate state of the outburst, which have been linked to the shock instability model of transonic accretion flow around black holes Chatterjee et al. (2024). Notably, Mondal et al. (2024) reported low polarization for this source, whereas Ling et al. (2024) found a null-detection.
@@
$\bullet$ IGR J17091-3624: This is a galactic BH-XRB transient system, discovered by INTEGRAL/IBIS in $2003$ (Kuulkers et al., 2003). Interestingly, IGR J17091362417091-3624 exhibits several exotic X-ray variability signatures, including the well known “heartbeat” class and therefore, often termed as the twin of GRS $1915+105$ (Capitanio et al., 2012; Radhika et al., 2018; Katoch et al., 2021; Wang et al., 2024; Shui et al., 2024). Using reflection spectroscopy, the inclination of the source is estimated as $\sim 30^{\circ}-45^{\circ}$ (Xu et al., 2017; Wang et al., 2018). In addition, spectral modeling suggests a black hole of mass 1012.310-12.3 $M_{\odot}$ likely to be present at the center of IGR J17091362417091-3624 (Iyer et al., 2015; Radhika et al., 2018). Based on the luminosity estimates, the distance of the source is predicted to lie within 111711-17 kpc (Rodriguez et al., 2011). Further, exceptionally high polarization degree has been reported for this source in LHS (Ewing et al., 2025; Debnath et al., 2025).
@@
$\bullet$ MAXI J1744-294: The sources is located near the Galactic center, is the most recently discovered X-ray transient by MAXI/GSC on January 2, 2025 at a flux level of $\sim 133$ mCrab (Kudo et al., 2025). The energy spectra of the source, obtained from follow-up NuSTAR observations, are found to be well described with the model comprising a thermal disc component of temperature $\sim 0.7$ keV, a power-law component of photon index $\sim 2.3$ and iron line profile at $\sim 6.6$ keV (Mandel et al., 2025). Based on these initial findings, the source is inferred to be a BH-XRB transient observed in the HSS (Mandel et al., 2025). Assuming a distance of $8$ kpc, the luminosity of the source is estimated to be $\sim 1.5\times 10^{37}$ erg $\rm s^{-1}$ in 2102-10 keV energy range (Mandel et al., 2025).
@@ 3 Data reduction
IXPE, NASA’s first space-based low-energy X-ray polarimetry mission, was launched on December 9, 2021 (Weisskopf et al., 2022). It comprises three identical gas-pixel detector units (DUs) designed to measure the polarimetric properties of astrophysical sources using the principle of photo-electron tracking in the 282-8 keV energy band. For each detected photon, the Stokes parameters (I, Q, U) are determined based on the azimuthal angle of the electric field vector, which is reconstructed by analyzing the photo-electron track within the detector plane. The additive nature of the Stokes parameters enables the summation of individual photon contributions to obtain the resultant I, Q, and U, which are used to calculate the polarization degree (PD) and polarization angle (PA).
@@
For the polarimetric analysis, we utilize the cleaned and calibrated level-2 event files from the three DUs of IXPE in the 282-8 keV energy range. Data reduction is performed using the latest IXPEOBSSIMv31.0.3 software (Baldini et al., 2022), following standard procedures outlined in Kislat et al. 2015; Strohmayer 2017; Kushwaha et al. 2023; Majumder et al. 2024a, b. The source and background regions are defined as a $60$ arcsec circular region centered at the source coordinates and an annular region with radii between $180$ and $240$ arcsec, respectively (Kushwaha et al., 2023; Majumder et al., 2024a). The XPSELECT task is used to extract source and background events from these regions. We use XPBIN task with various algorithms, such as PCUBE, PHA1, PHA1Q and PHA1U, to generate the necessary data products for model-dependent polarization as well as spectro-polarimetric studies, respectively. The latest response files provided by the software team are used in the fitting of Stokes spectra. The IXPE light curves of $2000$ s time bin are extracted from the event files using XSELECT.
@@
NICER (Gendreau et al., 2016) is the X-ray mission onboard the International Space Station (ISS), sensitive in the soft X-ray energy band of 0.2120.2-12 keV and capable of excellent observations for in-depth spectro-temporal studies. The NICER/XTI data is reduced using the standard data extraction software (NICERDASv13) integrated within HEASOFT V6.34 with the appropriate calibration database. We use the tool nicerl2 with standard filtering criteria to generate the clean event files. The spectral products are generated using the task nicerl3-spect with background model 3c50 (Remillard et al., 2022). The NICER light curves of desired time resolutions are generated using nicerl3-lc task.
@@
We analyze the NuSTAR (Harrison et al., 2013) data using the dedicated software NUSTARDAS available in HEASOFT V6.34. We use the task nupipeline to generate the cleaned event files from both FPMA and FPMB instrument onboard NuSTAR. A circular region of $60$ arcsec radii at the source position and away from it is considered to extract the source and background spectra, instrument response and ancillary files, respectively, using the task nuproduct. All the NuSTAR spectra are grouped with $25$ counts per bin to obtain better statistics during the spectral fitting.
@@
AstroSat (Agrawal, 2006) data are extracted using the latest reduction software LaxpcSoftv3.4.4.The standard procedures of AstroSat/LAXPC data extraction are performed following Antia et al. 2017, 2021; Antia et al. 2022; Sreehari et al. 2019b, 2020; Majumder et al. 2022. We use LAXPC20 data for generating the light curves in different energy ranges for the respective sources except Swift J1727.816131727.8-1613, for which LAXPC10 data in low gain mode was available for analysis (see Nandi et al., 2024, for details). Note that, level-1 data corresponding to single events from all layers of the detector are considered in each observation.
@@
We utilize on-demand [^3] http://maxi.riken.jp/mxondem/ processed MAXI/GSC (Matsuoka et al., 2009) data and scaled map transient analysis data [^4] https://swift.gsfc.nasa.gov/results/transients/ of Swift/BAT (Krimm et al., 2013) for the individual sources in different energy ranges. Note that, MAXI/GSC count rates in 2202-20 keV energy range are converted into the flux unit (Crab) following the relation $1$ Crab =3.8=3.8 photons $\rm cm^{-2}$ $\rm s^{-1}$ as suggested by the instrument team [^5] http://maxi.riken.jp/top/readme.html. Similarly, for Swift/BAT data in the 155015-50 keV energy range, we obtain the corresponding photon flux [^6] https://swift.gsfc.nasa.gov/results/transients/ in units of mCrab, where $1$ mCrab =0.000220=0.000220 counts $\rm cm^{-2}$ $\rm s^{-1}$.
@@ 4.1 Temporal Analysis
Figure 2: Results of temporal (top panel) and spectral (bottom panel) analyses of ten BH-XRBs in different spectral states (color coded) obtained using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations. Top panels: PDS obtained from NICER (0.5100.5-10 keV) and/or NuSTAR/AstroSat observations in different energy ranges are depicted for the respective sources in various epochs. The power spectrum of GX 3394339-4 from NICER observation is rescaled by a factor of $5$. Bottom panels: Energy spectra of the sources from NICER and/or NuSTAR data are shown during the respective epochs. PDS and energy spectra in LHS, HIMS, SIMS, HSS and SPL states are presented using red, blue, magenta, orange and green colors, respectively. PDS from NuSTAR and AstroSat data are plotted using circle and triangle, irrespective of the spectral states. See the text for details.
@@ 4.1.1 Multi-mission Monitoring
We study the variability properties of each source using multi-mission data during the epochs of IXPE observations. In Fig. 1, we present the count rate obtained with MAXI/GSC (2202-20 keV) and Swift/BAT (155015-50 keV) monitoring in panels (a) and (b) along with the variation of hardness ratio ($HR$) from MAXI/GSC in panel (c), respectively, for all sources under consideration. Here, $HR$ is defined as the flux ration in 6206-20 KeV and 262-6 keV. Note that, flux values associated with the count rates in soft (hard) energy bands with MAXI/GSC (Swift/BAT) are also mentioned in units of mCrab in the $y$-axis (right side) of the respective panels. Further, radio flux densities obtained from different observational campaigns, such as RATAN (Veledina et al., 2024a; Ingram et al., 2024), E-MERLIN (Williams-Baldwin et al., 2023), VLITE (Peters et al., 2023) and AMI-LA (Steiner et al., 2024) near the IXPE epochs, are shown in panel (d) for the respective sources based on their availability. We observe that all sources show marginal variation in both MAXI/GSC and Swift/BAT count rates over the entire IXPE exposure for a given epoch. Hence, in Fig. 1a-c, we present the average values of the respective quantities over the entire IXPE exposure of each epoch.
@@
Interestingly, for Swift J1727.816131727.8-1613, we observe that both MAXI/GSC and Swift/BAT flux decrease as its outburst progresses. The $HR$ also sharply decreases and anti-correlates with the accompanied radio flux density (see Fig. 1c-d). The variation of both low and high energy fluxes along with $HR$ in presence of radio ejection suggest that the source evolves through spectral state transitions. In particular, we notice exceptionally bright ($\gtrsim 3$ Crab) intermediate states ($HR\sim 0.1-0.5$, E1-E5) followed by relatively faint ($\sim 30$ mCrab) softer states ($HR\sim 0.04-0.1$, E6-E7) before the appearance of dimmed ($\sim 100$ mCrab) hard state ($HR\sim 0.3$, E8) (see Table 2). Moreover, these variability features resemble the ‘canonical’ state transition as HIMS $\rightarrow$ SIMS $\rightarrow$ HSS $\rightarrow$ LHS (decay), commonly exhibited by BH-XRB transients (see Nandi et al., 2012, for details). Similarly, for Cyg X1-1, we observe a marginal increase in the MAXI/GSC count rate and a corresponding decrease in the Swift/BAT count rate during epochs E3 to E7. In the subsequent observations (E8 to E11), the MAXI/GSC count rate decreases while the Swift/BAT count rate rises. In addition, $HR$ shows a significant drop from $0.4$ to $0.1$ followed by an increase in values exceeding $0.4$. This trend possibly indicates the spectral state transitions between LHS (E1-E2, E8-E11) and HSS (E3-E7, E12), respectively (see Table 1).
@@
However, only marginal variations are seen in both low- and high-energy count rates for several sources, such as LMC X1-1, LMC X3-3, 4U $1957+115$, GX 3394339-4, Swift J151857.0572147151857.0-572147 and MAXI J17442941744-294. For all of them, the hardness ratio remains low ($HR\lesssim 0.1$) and these sources are also detected at flux levels below approximately $100$ mCrab. These possibly suggest that the sources are likely in softer spectral states. In contrast, Cyg X3-3 exhibits a clear sequence of state transitions, evolving from the LHS to the HSS, then to the SIMS, and back to HSS. This evolution is accompanied by moderate changes in count rates and hardness ratio, and is likely associated with strong radio flares ($\sim 100$ mJy) (see Table 1, Fig. 1, and Veledina et al. 2024a). For 4U 1630471630-47, HSS (Kushwaha et al., 2023) and SPL (Rodriguez Cavero et al., 2023) states are observed with reasonable variations in the count rates and $HR$. It is worth noting that IGR 17091362417091-3624 is observed in the LHS at $\sim 1$ Crab MAXI/GSC flux with $HR\sim 0.4$, very similar to the LHS observations of Cyg X1-1. Furthermore, GRS $1915+105$ is observed with IXPE in its present faint and obscured state with a flux level of $\sim 6$ mCrab only (see Fig. 1). We find that observations in harder spectral states (LHS and HIMS) and relatively softer spectral states (HSS, SIMS and SPL) of the sources are separated in two distinct regions of $HR$ variations (see Fig. 1). In particular, $HR\sim 0.22$ separates two regions, with $HR>0.22$ indicating harder and $HR<0.22$ corresponds to softer spectral states with few exceptions.